Richard has been given a 6 question multiple choice quiz in hi history clad. Each question has f e answer of which only one is correct. Since Richard has not attended the class recently he doesn't know any of the answers. Assuming that Richard guesses on all 6 questions find the probability that he will answer all the questions right. Round your answer to nearest hundredth
Answer by stanbon(75887)

Richard has been given a 6 question multiple choice quiz in history class. Each question has five answers of which only one is correct. Since Richard has not attended the class recently he doesn't know any of the answers. Assuming that Richard guesses on all 6 questions find the probability that he will answer all the questions right. Round your answer to nearest hundredth
----
Binomial Problem with n = 6 and p(correct) = 1/5
----
P(x = 6) = (1/5)^6 = 0.000064
======================================
Cheers,
Stan H.
===========